# LeapFrog: Getting the Jump on Multi-Scale Materials Simulations Using Machine Learning

### v1

@@ Abstract
The development of novel materials in recent years has been accelerated greatly by the use of computational modelling techniques aimed at elucidating the complex physics controlling microstructure formation in materials, the properties of which control material function. One such technique is the phase field method, a field theoretic approach that couples various thermophysical fields to microscopic order parameter fields that track the phases of microstructure. Phase field models are framed as multiple, non-linear, partial differential equations, which are extremely challenging to compute efficiently. Recent years have seen an explosion of computational algorithms aimed at enhancing the efficiency of phase field simulations. One such technique, adaptive mesh refinement (AMR), dynamically adapts numerical meshes to be highly refined around steep spatial gradients of the PDE fields and coarser where the fields are smooth. This reduces the number of computations per time step significantly, thus reducing the total time of computation. What AMR doesn’t do is allow for adaptive time stepping. This work combines AMR with a neural network algorithm that uses a U-Net with a Convolutional Long-Short Term Memory (CLSTM) base to accelerate phase field simulations. Our neural network algorithm is described in detail and tested in on simulations of directional solidification of a dilute binary alloy, a paradigm that is highly practical for its relevance to the solidification of alloys.
@@ 2.1 Phase Field Model
Figure 1: Selection of snapshots from different times in the growth of a dendrite array simulated from Eq. (3) and initiated form the morphology in the left frame.
@@ 2.2.2 Loss Functions
Figure 4: Taking the difference between the system’s fields at two different times (first two columns) to construct the left-hand sides of Eq. 4 & 5 (third column). It is clear that values of 0 occupy the majority of the $\Delta\{\phi,C\}(\mathbf{x},t)$ fields (blue in the case of $\phi$, and orange in the case of $C$). The final column results from the construction of a binary mask after the application of a custom re-scaling function to emphasize the interface and solute diffusion.
@@ 3.1 Acceleration
Figure 5: Comparison of the time savings and final results between traditional Adaptive Mesh Refinement (AMR) and neural network accelerated AMR phase field simulations, with $N_{ML}/N_{PF}=5/2$.
@@
Table 1: Comparison of wall times required per $1000dt$ of simulation time. When not indicated otherwise, the values listed are purely the time from the beginning of the computation of $1000dt$ to the end, not taking into consideration overhead (the time required to load up the system data and write the results to a file). The overhead required by the LeapFrog’s current [proof-of-concept] implementation is the majority of its wall time cost.
@@ 3.2 Adaptive Time Stepping
Figure 6: Comparison of the time savings and final results between traditional Adaptive Mesh Refinement (AMR) and neural network accelerated AMR phase field simulations, with $N_{ML}/N_{PF}=1$.
@@ 3.3 Prediction Quality of $\phi$ and $C$ Fields
Figure 7: Comparison of concentration profiles along a dendrite’s core (centerline) for two simulations, one evolved with Adaptive Mesh Refinement (AMR) (right frame, dashed line) and the other with the LeapFrog (LF) algorithm (right frame, continuous line). The left frame is split into two sub-frames, with the left sub-frame showing the centerline along which $C$ is measured from the AMR output, and the right sub-frame showing that of from the LeapFrog outputs. In the right frame. The point $y=0$ corresponds to the top of the centerline in the left sub-frames.
@@
Figure 8: Generation of mask used to smooth dendrite concentration profiles output from the LeapFrog algorithm (a), by taking the magnitude of the $C$ field’s gradients (b), constraining it to the solid phase bulk as well as a to threshold magnitude (c), and zeroing-out the portion of the system generated during the initial transient phase of the simulation (anywhere with $y<V_{p}\frac{65000dt}{dx}$).
@@
Figure 9: Comparison of concentration profiles along a dendrite’s core (centerline) for a simulation evolved with Adaptive Mesh Refinement (AMR) and LeapFrog (LF) (red lines in Fig. 7) with smoothing mask illustrated in Fig. 8 applied to the $C$-field output from the LeapFrog algorithm. The core (centerline) concentration profile is shown on the left. Also shown are the concentration profiles along the two isotherms in the liquid (top right) and solid (bottom right) phases. The point $y=0$ corresponds to the top of the centerline in the left sub-frames.
@@
Figure 10: Mean power spectra comparison with uncertainty intervals between direct Phase F“ield(PF)/AMR simulations (black) and LeapFrog(LF)-enabled PF simulation (red) for both the phase (left) and concentration (right) fields. Red/Black barred areas designate regions of overlap between the PF and LF uncertainty. The power spectra have been constructed from 10 simulations with different randomly initialized seeds, and the $k=0$ mode set to 0.

### v2

# LeapFrog: Accelerating Multiscale Materials Simulations with Machine Learning

@@ Abstract
Recent developments of novel materials have been greatly accelerated by computational modelling techniques that elucidate the complex physics controlling microstructure formation, the properties of which determine material function. The phase field (PF) method evolves said microstructure phases by coupling thermophysical and microscopic order parameter fields through multiple non-linear and costly to compute partial differential equations. Adaptive mesh refinement (AMR) significantly reduces the number of computations per time step, and thus the total computation time, by dynamically adapting numerical meshes resolution in proportion to local gradients. What AMR doesn’t do is allow for adaptive time stepping. This work combines AMR with a neural network algorithm that uses a U-Net with a Convolutional Long-Short Term Memory (CLSTM) base to accelerate PF simulations. Our neural network algorithm is described in detail and tested on directional dilute binary alloy solidification simulations, a highly practical paradigm in alloy solidification.
@@ 2.1 Phase Field Model
Figure 1: Selection of snapshots from different times in the growth of a dendrite array.
Simulated from Eq. (3) and initiated from the morphology in the left frame.
@@ 2.2.2 Loss Functions
Figure 4: Interface & Gradient mask generation. Taking the difference between the system’s fields at two different times (first two columns) to construct the left-hand sides of Eq. 4 & 5 (third column). It is clear that values of 0 occupy the majority of the $\Delta\{\phi,C\}(\mathbf{x},t)$ fields (blue in the case of $\phi$, and orange in the case of $C$). The final column results from the construction of a binary mask after the application of a custom re-scaling function to emphasize the interface and solute diffusion.
@@ 3.1 Acceleration
Figure 5: 5:2 Leapfrog Algorithm Results. Comparison of the time savings and final results between traditional Adaptive Mesh Refinement (AMR) and neural network accelerated AMR phase field simulations, with $N_{ML}/N_{PF}=5/2$.
@@
Table 1: LeapFrog Performance Metrics. Comparison of wall times required per $1000dt$ of simulation time. When not indicated otherwise, the values listed are purely the time from the beginning of the computation of $1000dt$ to the end, not taking into consideration overhead (the time required to load up the system data and write the results to a file). The overhead required by the LeapFrog’s current [proof-of-concept] implementation is the majority of its wall time cost.
@@ 3.2 Adaptive Time Stepping
Figure 6: 5:5 LeapFrog Algorithm results. Comparison of the time savings and final results between traditional Adaptive Mesh Refinement (AMR) and neural network accelerated AMR phase field simulations, with $N_{ML}/N_{PF}=1$.
@@ 3.3 Prediction Quality of $\phi$ and $C$ Fields
Figure 7: PF and LeapFrog Concentration Profiles Comparison. Comparison of concentration profiles along a dendrite’s core (centerline) for two simulations, one evolved with Adaptive Mesh Refinement (AMR) (right frame, dashed line) and the other with the LeapFrog (LF) algorithm (right frame, continuous line). The left frame is split into two sub-frames, with the left sub-frame showing the centerline along which $C$ is measured from the AMR output, and the right sub-frame showing that of from the LeapFrog outputs. In the right frame. The point $y=0$ corresponds to the top of the centerline in the left sub-frames.
@@
Figure 8: Generation of mask used to smooth dendrite concentration profiles output from the LeapFrog algorithm
a) The concentration field of a system reaching the end of the dendritic growth regime.
b) The magnitude of the gradients in said concentration field.
c) Constraining the mask to areas where $\phi$ corresponds to the solid phase as well filtering the concentration gradient magnitude through a magnitude threshold.
d) Restricting the resulting mask to the portion of the system generated after the initial transient phase of the simulation ($y>V_{p}\frac{65000dt}{dx}$).
@@
Figure 9: Comparison of concentration profiles along a dendrite’s core (centerline) for a simulation evolved with Adaptive Mesh Refinement (AMR) and LeapFrog (LF) at $t^{*}=250000dt$. (red lines in Fig. 7) with smoothing mask illustrated in Fig. 8 applied to the $C$-field output from the LeapFrog algorithm. The core (centerline) concentration profile is shown on the left. Also shown are the concentration profiles along the two isotherms in the liquid (top right) and solid (bottom right) phases. The point $y=0$ corresponds to the top of the centerline in the left sub-frames.
@@
Figure 10: Mean power spectra comparison with uncertainty intervals between direct Phase Field(PF)/AMR simulations (black) and LeapFrog(LF)-enabled PF simulation (red) for both the phase (left) and concentration (right) fields. Red/Black barred areas designate regions of overlap between the PF and LF uncertainty. The power spectra have been constructed from 10 simulations with different randomly initialized seeds, and the $k=0$ mode set to 0.
@@ 4 Discussion
Acknowledgements
We acknowledge support from the Mitacs Organization through its Accelerate Fellowship program and Hydro-Québec. We also acknowledge the computational resources made available by the Digital Research Alliance of Canada.
Author Contributions
D.P. designed and trained the ML model, generated the data for model training and validation, implemented the LeapFrog algorithm, performed analyses of the results and wrote the paper. M.G. provided PF AMR code, found simulation parameters that demonstrated significant side-branching. M.G. and N.P advised the ML model and LeapFrog algorithm design. N.P. supervised all work, and edited the paper.
Competing Interests
The authors declare no competing financial or non-financial interests.
Data Availability
The datasets used and/or analyzed during the current study available from the corresponding author on reasonable request.
Code Availability
The underlying code for this study and training/validation datasets are not publicly available, but may be made available to qualified researchers on reasonable request from the corresponding author.
Additional Information
Correspondence and requests for materials should be addressed to Damien Pinto or Nikolas Provatas.
Reprints and permission information is available at nature.com/nature-portfolio/reprints-and-permissions.